Given that 43+x/810x = 210-2x/643x what is the value of x?


1
Expert's answer
2021-08-22T17:25:46-0400

Let us solve the equation


"\frac{4^{3+x}}{8^{10x}} = \frac{2^{10-2x}}{64^{3x}}"


which is equivalent to the following equations:


"\frac{2^{2(3+x)}}{2^{30x}} = \frac{2^{10-2x}}{2^{18x}}"


"2^{6+2x-30x} = 2^{10-2x-18x}"


"2^{6-28x} = 2^{10-20x}"


"6-28x= 10-20x"


"8x= -4"


"x=-\frac{1}{2}"


We conclude that the value of "x" is "-\frac{1}{2}."